Solve Problems Using Linear Systems
2. James looks in his TV cabinet and finds some old Beta and VHS tapes. He has 17 tapes in all. He finds that he has three more Beta tapes than VHS tapes. How much many of each type does he have?
(The answer is: 10 Beta tapes and 7 VHS tapes, but i dont know how solve this problem?) Thank you very much!!!

Answer by scott8148(6628)

"He has 17 tapes in all" ___ B+V=17

"He finds that he has three more Beta tapes than VHS tapes" ___ B=V+3

substituting ___ (V+3)+V=17 ___ 2V+3=17 ___ 2V=14 ___ V=7

substituting ___ B+(7)=17 ___ B=10